# FDSG: Forecasting Dynamic Scene Graphs

### v1

@@ 3.6 Training and Inference
where $\mathcal{L}_{\text{match}}(r(k),\hat{r}(\rho(k)))$ is the matching loss to assign predicted triplet $\hat{r}$ with index $\rho(i)$ to ground truth triplet $r(i)$. This matching loss is a sum of entity classification loss, predicate classification loss and entity bounding box regression loss:
@@
where $\{s,o,p\}$ denotes subject, object, predicate respectively, and $\alpha_{l}$ and $\beta_{l}$ are weighting factors. We use focal loss [36] for classification $\mathcal{L}_{\text{cls}}$, and a weighted sum of L1-loss and GIoU-loss for bounding box regression $\mathcal{L}_{\text{box}}$. When calculating the predicate classification loss, we follow OED [41] and ignore the loss between the prediction and any padded ground truth.
We use the same Hungarian Matching loss for SGF. During training, the model only outputs forecasted triplets for the target frame. We select a reference frame and use it to forecast the scene graph for the target frame, as specified in Eq. (10). Therefore, both the Scene Graph Generation Module and the Scene Graph Forecast Module output predictions for the same target frame. Unlike SGA, we do not need to store additional samples or labels for multiple observed and anticipatory targets.
@@
For SGA and SGF, our model only sees part of the video $\{V_{1},...,V_{T}\}$ and samples $n$ reference frames $\{V_{\text{ref}1},...,V_{\text{ref}n}\}$ from the observed part. Our model inputs in this case are $\{V_{\text{ref}1},...,V_{\text{ref}n}\}$ and the last observed frame $V_{T}$. Our model generates a scene graph $\hat{G}_{T}$ for the last observed frame, and from $\hat{G}_{T}$ it forecasts scene graphs $\{\hat{G}_{t}\}_{t=T+1}^{H}$ for a future time window of length $H-T$.
4 Experiments
Table 1: Comparison with state-of-the-art Dynamic Scene Graph Generation (DSGG) methods on Action Genome for Scene Graph Detection (SGDET). Best values are highlighted in boldface.
@@ 4 Experiments
Ours 35.3 42.9 49.8 37.2 47.2 56.5 22.2 27.8 33.0 27.8 42.0 54.1
Table 2: Comparison with state-of-the-art Scene Graph Anticipation (SGA) methods in the Action Genome Scenes (AGS). Best values are highlighted in boldface.
AGS, $\mathcal{F}=0.5,IoU\geq 0$ Recall@K (R@K) $\uparrow$ Mean Recall@K (mR@K) $\uparrow$
@@
STTran++ [9, 33] 19.7 30.2 31.8 16.6 29.1 51.5 6.3 11.3 12.3 6.6 14.7 33.4
@@
Ours 28.3 36.5 45.3 30.1 40.4 52.2 18.1 23.5 27.5 23.2 34.3 45.8
4.1 Experimental Settings
Dataset. We evaluate the proposed model on the Action Genome dataset [16], which offers frame-level annotations for dynamic scene graph generation tasks. It encompasses 234,253 frames with 476,229 bounding boxes representing 35 object classes (excluding “person”) and 1,715,568 instances across 25 relationship (predicate) classes. These relationships are categorized into three types: attentional (e.g., a person looking at an object), spatial (e.g., object positions relative to each other), and contacting (e.g., interactions such as holding or eating). Notably, this dataset supports multi-label annotations, with 135,484 subject-object pairs annotated with multiple spatial or contact relationships, facilitating nuanced analysis of complex human-object interactions. This allows us to follow the two widely recognized strategies [3, 22] to build scene graphs and evaluate their performance: (1) With Constraint: enforcing a unique interaction for each pair. (2) No Constraint: allowing multiple relationships between the same pair.
Evaluation Metrics. We adopt the standard metrics Recall@K (R@K) and mean Recall@K (mR@K) to evaluate our model’s performance under both With Constraint and No Constraint strategies, where higher values indicate better coverage of true relationships. Following previous works, we set $K\in\{10,20,50\}$. For DSGG and SGF, we primarily evaluate our model using the most challenging protocol, Scene Graph Detection (SGDET), which requires predicting entity labels, bounding boxes, and pairwise relationships for a target frame. A predicted entity bounding box is considered correct if its Intersection over Union (IoU) with the ground truth box exceeds 0.5. Additional details on DSGG evaluation protocols Scene Graph Classification (SGCLS) and Predicate Classification (PredCLS) are provided in the Supplementary Materials. For SGA and SGF, to split the observation and forecasting time horizons, we use $\mathcal{F}$, the fraction of observed frames within the total video frames. A typical value is $\mathcal{F}=0.5$, meaning the first half of the video is observed, while the second half is used for evaluation. For SGA, our model is evaluated on Action Genome Scenes, Partially Grounded Action Genome Scenes, and Grounded Action Genome Scenes, following [33]. Notably, in SGA, bounding box is not evaluated [33], therefore we use an IoU threshold of 0.
Implementation Details. Our model is initialized with 4-scale DINO pretrained for 36 epochs on the COCO 2017 Object Detection Dataset [25], which consists of 6 Transformer layers for encoder and decoder respectively, and uses ResNet50 [14] as the backbone. The number of entity queries is set to 100. Following OED [41], we set $\text{TopK}=\{80n,50n,30n\}$, where $n=2$, denoting the number of reference frames. For the NeuralSDE in our Triplet Dynamics Model, we employ a solver with a reversible Heun method, the same as in SceneSayer [33]. $\mathcal{F}$ is set to $\{0.3,0.5,0.7,0.9\}$. More details of our model training are provided in Supplementary B.4.
Table 3: The performance of Scene Graph Forecasting (SGF) on Action Genome for Scene Graph Detection (SGDET). Best values are highlighted in boldface.
SGDET, $\mathcal{F}=0.5$ Recall@K (R@K) $\uparrow$ Mean Recall@K (mR@K) $\uparrow$
$IoU\geq 0.5$ With Constraint No Constraint With Constraint No Constraint
Method 10 20 50 10 20 50 10 20 50 10 20 50
SceneSayerODE+ [33] 0.9 1.0 1.1 0.9 1.2 1.6 0.5 0.6 0.6 0.6 1.0 1.7
SceneSayerSDE+ [33] 2.5 3.2 3.4 2.6 3.7 5.0 1.1 1.6 1.7 1.4 2.2 4.4
Ours 10.6 13.3 15.5 12.0 15.5 18.8 8.4 11.2 12.6 11.4 17.2 22.5
4.2 Quantitative Results and Comparison
@@
Dynamic Scene Graph Generation. We evaluate our method on DSGG using the SGDET protocol, with results summarized in Table 1. Our approach consistently outperforms state-of-the-art methods across all metrics in both With Constraint and No Constraint settings, in particular achieving a notable performance gain over OED [41] in both R@50 and mR@50. For comparison in the PredCLS and SGCLS tasks, please refer to Supplementary C.1.
Scene Graph Anticipation. Table 2 presents our results on SGA with an observation fraction of $\mathcal{F}=0.5$ on Action Genome Scenes. Our method outperforms SceneSayer [33] across all metrics, achieving up to 50% improvement in mR@10 and mR@50, and averaging a 40% gain across Mean Recall metrics, highlighting our model’s effectiveness in reducing prediction bias. Additional results for $\mathcal{F}={0.3,0.7,0.9}$ and for Grounded / Partially Grounded protocols are provided in Supplementary C.2.
Scene Graph Forecasting. We report our SGF results at $\mathcal{F}=0.5$ in Table 3. Baselines are adapted from SceneSayerODE/SDE [33], where predicted predicate representations are fed into MLPs for bounding box forecasting (see Supplementary B.5 for details). While these models perform well for predicate anticipation, they show limited capacity for entity forecasting. In contrast, our model significantly outperforms them, demonstrating a stronger ability to learn temporal dynamics. Together with our improvements in DSGG and SGA (Tables 1, 2), this highlights the advantages of our forecasting-driven design.
Entity Forecasting in SGF. We explicitly evaluate entity forecasting using Object Detection Recall. At $\mathcal{F}=0.5$ and IoU $\geq 0.5$, FDSG achieves a recall of 43.85, significantly outperforming a baseline recall of 40.33 that reuses entities from the last observed frame. See Supplementary C.3 for detailed evaluation with different $\mathcal{F}$ and IoU thresholds.

### v2

@@ 3.6 Training and Inference
where $\mathcal{L}_{\text{match}}(r(k),\hat{r}(\rho(k)))$ is the matching loss to assign predicted triplet $\hat{r}$ with index $\rho(i)$ to ground truth triplet $r(i)$. This matching loss is a sum of entity classification loss, predicate classification loss, and entity bounding box regression loss:
@@
where $\{s,o,p\}$ denotes subject, object, and predicate, respectively, and $\alpha_{l}$ and $\beta_{l}$ are weighting coefficients. We use focal loss [41] for classification $\mathcal{L}_{\text{cls}}$, and a weighted sum of L1-loss and GIoU-loss for bounding box regression $\mathcal{L}_{\text{box}}$. When calculating the predicate classification loss, we follow OED [11] and ignore the loss between the prediction and any padded ground truth.
For the SGF task, we adopt the same Hungarian Matching loss. During training, the model generates forecasted triplets only for the target frame. A reference frame is selected to guide the prediction of the scene graph for the target frame, as defined in Eq. (16). Consequently, both the Scene Graph Generation Module and the Forecast Module produce predictions for the same target frame. This is different from the SceneSayer training scheme for SGA [15], which requires storing all observed samples and multiple anticipation targets.
@@
For SGA and SGF, our model only sees part of the video $\{V_{1},...,V_{T}\}$ and samples $n$ reference frames $\{V_{\text{ref}1},...,V_{\text{ref}n}\}$ from the observed part. The inputs of our model in this case are $\{V_{\text{ref}1},...,V_{\text{ref}n}\}$ and the last observed frame $V_{T}$. Our model generates a scene graph $\hat{G}_{T}$ for the last observed frame, and from $\hat{G}_{T}$ it forecasts scene graphs $\{\hat{G}_{t}\}_{t=T+1}^{H}$ for a future time window of length $H-T$.
4 Evaluation for Different Tasks
TABLE II: Comparison with state-of-the-art Dynamic Scene Graph Generation (DSGG) methods on Action Genome for Scene Graph Detection (SGDET). Best values are highlighted in boldface.
@@ 4 Evaluation for Different Tasks
FDSG (ours) 35.3 42.9 49.8 37.2 47.2 56.5 22.2 27.8 33.0 27.8 42.0 54.1
We adopt the standard metrics Recall@K (R@K) and mean Recall@K (mR@K) to evaluate our model’s performance under both With Constraint and No Constraint strategies, where higher values indicate better coverage of true relationships. Following previous works, we set $K\in\{10,20,50\}$.
We summarize the evaluation of different tasks in Table I, and present detailed elaboration as follows.
4.1 Dynamic Sene Graph Generation (DSGG)
We primarily evaluate our model using the most challenging protocol, Scene Graph Detection (SGDET), which requires predicting entity labels, bounding boxes, and pairwise relationships for a target frame. A predicted entity bounding box is considered correct if its Intersection over Union (IoU) with the ground truth box exceeds 0.5. We also provide evaluation results for Predicate Classification (PredCLS) and Scene Graph Classification (SGCLS) for DSGG. PredCLS aims to classify the predicate labels with ground-truth bounding boxes and labels of the entities being given, while SGCLS aims to classify entity and predicate labels with only ground-truth bounding boxes being given.
TABLE III: Comparison with state-of-the-art Dynamic Scene Graph Generation methods on Action Genome for Scene Graph Classification (SGCLS). Best and second-best values are highlighted in boldface and underlined, respectively.
SGCLS Recall@K (R@K) $\uparrow$ Mean Recall@K (mR@K) $\uparrow$
@@ 4.1 Dynamic Sene Graph Generation (DSGG)
RelDN [42] 44.3 45.4 45.4 52.9 62.4 65.1 3.3 3.3 3.3 7.5 18.8 33.7
VCTree [43] 44.1 45.3 45.3 52.4 62.0 65.1 - - - - - -
TRACE [8] - 45.7 46.8 - - - 8.9 8.9 8.9 31.9 42.7 46.3
GPS-Net [21] 45.3 46.5 46.5 53.6 63.3 66.0 - - - - - -
STTran [6] 46.4 47.5 47.5 54.0 63.7 66.4 16.6 20.8 22.2 20.9 29.7 39.2
APT [14] 47.2 48.9 48.9 55.1 65.1 68.7 - - - - - -
DSG-DETR [29] 50.8 52.0 52.0 59.2 69.1 72.4 - - - - - -
TEMPURA [44] 47.2 48.3 48.3 56.3 64.7 67.9 30.4 35.2 35.2 48.3 61.1 66.4
FDSG (ours) 54.8 56.5 56.5 59.0 67.2 72.4 34.9 36.8 36.8 48.9 63.8 74.0
TABLE IV: Comparison with state-of-the-art Dynamic Scene Graph Generation methods on Action Genome for Predicate Classification (PredCLS). Best and second-best values are highlighted in boldface and underlined, respectively.
PredCLS Recall@K (R@K) $\uparrow$ Mean Recall@K (mR@K) $\uparrow$
With Constraint No Constraint With Constraint No Constraint
Method 10 20 50 10 20 50 10 20 50 10 20 50
RelDN [42] 66.3 69.5 69.5 75.7 93.0 99.0 6.2 6.2 6.2 31.2 63.1 75.5
VCTree [43] 66.0 69.3 69.3 75.5 92.9 99.3 - - - - - -
TRACE [8] 27.5 27.5 27.5 72.6 91.6 96.4 15.2 15.2 15.2 50.9 73.6 82.7
GPS-Net [21] 66.8 69.9 69.9 76.0 93.6 99.5 - - - - - -
STTran [6] 68.6 71.8 71.8 77.9 94.2 99.1 37.8 40.1 40.2 51.4 67.7 82.7
APT [14] 69.4 73.8 73.8 78.5 95.1 99.2 - - - - - -
TEMPURA [44] 68.8 71.5 71.5 80.4 94.2 99.4 42.9 46.3 46.3 61.5 85.1 98.0
OED [11] 73.0 76.1 76.1 83.3 95.3 99.2 - - - - - -
FDSG (ours) 72.9 75.9 75.9 83.2 95.4 99.4 41.0 44.7 44.8 56.7 83.3 97.7
4.2 Scene Graph Anticipation (SGA)
We use $\mathcal{F}$, the fraction of observed frames within the total video frames, to split the observation and forecasting time horizons. A typical value is $\mathcal{F}=0.5$, meaning the first half of the video is observed, while the second half is used for evaluation. For SGA, our model is evaluated on Action Genome Scenes (AGS), Partially Grounded Action Genome Scenes (PGAGS), and Grounded Action Genome Scenes (GAGS), following [15]. AGS resembles SGDET, where the model only sees video frames and predicts future scene graphs without bounding boxes. Similarly, GAGS aims to anticipate predicate labels, given ground-truth bounding boxes and labels of the entities. PGAGS aims to classify entity and predicate labels with ground-truth bounding boxes given.
4.3 Scene Graph Forecasting (SGF)
As the task of SGF is to forecast complete scene graphs, we adopt the SGDET to evaluate the prediction of entity labels, bounding boxes, and predicates for an unobserved frame in the future. SGF is evaluated under different observed fractions $\mathcal{F}$. A notable difference between SGDET for SGF and AGS for SGA is that SGDET evaluates bounding box predictions with an IoU threshold of 0.5, while bounding box is not evaluated in AGS, which is equivalent to using an IoU threshold of 0. In addition, we supplement SGF assessment by evaluating entity forecasting with object detection recall and average precision.
5 Experiments
5.1 Experimental Settings
5.1.1 Dataset
We evaluate the proposed model on the Action Genome dataset [18], which offers frame-level annotations for dynamic scene graph generation tasks. It encompasses 234,253 frames with 476,229 bounding boxes representing 35 object classes (excluding “person”) and 1,715,568 instances across 25 relationship (predicate) classes. These relationships are categorized into three types: attentional (e.g., a person looking at an object), spatial (e.g., object positions relative to each other), and contacting (e.g., interactions such as holding or eating). Notably, this dataset supports multi-label annotations, with 135,484 subject-object pairs annotated with multiple spatial or contact relationships, facilitating nuanced analysis of complex human-object interactions. This allows us to follow the two widely recognized strategies [45, 46] to build scene graphs and evaluate their performance: (1) With Constraint: enforcing a unique interaction for each pair. (2) No Constraint: allowing multiple relationships between the same pair.
5.1.2 Implementation Details
Our model is initialized with a 4-scale DINO pretrained for 36 epochs on the COCO 2017 Object Detection Dataset [47], which consists of 6 Transformer layers for the encoder and decoder, respectively, and uses ResNet50 [48] as the backbone. The number of entity queries is set to 100. Each entity content representation has 256d and each triplet representation $Z$ has 512d, resulting in $Z_{T_{g}}\in\mathbb{R}^{100\times 512}$ in Eq. (16). Following OED [11], we set $\text{TopK}=\{80n,50n,30n\}$, where $n=2$, denoting the number of reference frames. For the NeuralSDE in our Triplet Dynamics Model, we employ a solver with a reversible Heun method, the same as in SceneSayer [15]. $\mathcal{F}$ is set to $\{0.3,0.5,0.7,0.9\}$.
5.1.3 Model Training
We first pre-trained the Scene Graph Generation baseline, as in OED [11], and then fine-tuned the Temporal Aggregation Module together with the Forecast Module. The Scene Graph Generation baseline was trained for 5 epochs with a starting learning rate of 0.0001 and was reduced by a factor of 0.1 at the 4th epoch. We used AdamW [49] as optimizer with a batch size of 2. The temporal and forecast modules were fine-tuned for 3 epochs, with a starting learning rate of 0.0001, batch size of 1, and reduced learning rate by a factor of 0.1 at the 3rd epoch. The weighting coefficients $\alpha_{l}$ and $\beta_{l}$ in the matching loss (Eq. (19)) are $1.0$ and $5.0$, respectively.
5.2 Quantitative Results and Comparison
5.2.1 Dynamic Scene Graph Generation
TABLE V: Comparison with state-of-the-art Scene Graph Anticipation methods on Action Genome Scenes (AGS). Best and second-best values are highlighted in boldface and underlined, respectively.
With Constraint No Constraint With Constraint No Constraint
10 20 50 10 20 50 10 20 50 10 20 50
FDSG (ours) 26.8 34.2 43.0 27.8 37.6 48.7 15.4 20.2 24.7 18.3 29.8 41.7
@@ 5.2.1 Dynamic Scene Graph Generation
FDSG (ours) 28.3 36.5 45.3 30.1 40.4 52.2 18.1 23.5 27.5 23.2 34.3 45.8
FDSG (ours) 35.5 43.9 50.6 37.1 49.2 60.2 21.6 27.9 31.2 25.0 41.9 54.7
FDSG (ours) 43.8 52.8 61.1 47.5 59.2 70.9 24.1 30.1 34.9 30.9 44.6 59.5
TABLE VI: Comparison with state-of-the-art Scene Graph Anticipation methods on Partially Grounded Action Genome Scenes (PGAGS). Best and second-best values are highlighted in boldface and underlined, respectively.
With Constraint No Constraint With Constraint No Constraint
10 20 50 10 20 50 10 20 50 10 20 50
FDSG (ours) 31.3 32.4 32.4 35.2 41.8 47.1 20.6 21.8 21.8 28.9 40.8 51.4
FDSG (ours) 34.3 35.2 35.2 38.1 45.0 50.5 22.0 23.4 23.4 31.0 43.0 53.7
FDSG (ours) 39.9 40.7 40.7 44.3 51.5 57.3 24.7 25.9 25.9 33.8 47.7 58.1
FDSG (ours) 45.9 46.6 46.6 53.0 60.0 64.7 27.9 28.8 28.8 41.8 54.0 66.6
TABLE VII: Comparison with state-of-the-art Scene Graph Anticipation methods on Grounded Action Genome Scenes (GAGS). Best and second-best values are highlighted in boldface and underlined, respectively.
With Constraint No Constraint With Constraint No Constraint
10 20 50 10 20 50 10 20 50 10 20 50
FDSG (ours) 46.2 48.5 48.5 55.0 67.3 73.1 24.6 27.6 27.7 31.3 51.6 73.8
FDSG (ours) 51.1 53.2 53.2 60.9 73.8 79.4 26.2 29.3 29.3 34.4 56.0 77.3
FDSG (ours) 57.3 59.1 59.1 69.0 81.8 87.1 29.2 32.2 32.2 39.0 62.9 85.9
FDSG (ours) 63.8 65.1 65.1 79.0 90.4 94.5 33.3 35.3 35.3 48.7 72.0 92.2
@@
Our evaluation results on DSGG using the SGDET protocol are summarized in Table II. Our approach consistently outperforms state-of-the-art methods across all metrics in both With Constraint and No Constraint settings, in particular achieving a notable performance gain over OED [11] in both R@50 and mR@50.
Further evaluation results for Scene Graph Classification (SGCLS) and Predicate Classification (PredCLS) for the DSGG task are presented in Table III and Table IV, respectively. It is worth mentioning that, unlike OED [11], we do not train separate models for the PredCLS and SGCLS tasks. We follow the recently verified match-and-assign strategy in one-stage scene graph generation methods [27, 28], the ground-truth labels can be easily assigned to the queries with matched triplet proposals for the evaluation of PredCLS and SGCLS. In contrast, OED follows the two-stage model strategy which the model is modified to incorporate additional input information (ground truth bounding boxes in SGCLS and additionally ground truth bounding boxes and entity labels in PredCLS) and then re-trained.
For SGCLS, in general, our model outperforms other baselines, especially TEMPURA [44], which is specifically designed for unbiased scene graph generation on the Mean Recall metrics. Since the SGCLS results of OED are not reported in the paper [11] and the model checkpoint is unavailable, we have to exclude OED in Table III. For PredCLS, our model achieves comparable performance to the state-of-the-art model OED, even without training a separate model for PredCLS.
The strong performances of our FDSG method on all three protocols are consistent, demonstrating the overall superiority of FDSG on the DSGG task.
5.2.2 Scene Graph Anticipation
TABLE VIII: Comparison with baseline methods on Action Genome for the Scene Graph Forecasting (SGF) task. Best values are highlighted in boldface.
With Constraint No Constraint With Constraint No Constraint
10 20 50 10 20 50 10 20 50 10 20 50
0.3 SceneSayerODE+ [15] 3.6 4.3 4.4 3.8 4.6 5.3 2.4 2.9 2.9 2.8 3.2 3.9
SceneSayerSDE+ [15] 4.3 4.7 5.1 4.5 5.3 6.2 2.5 3.5 3.5 2.9 4.0 4.7
FDSG (ours) 9.2 11.4 13.1 10.2 13.1 16.2 7.6 9.7 11.3 9.2 14.2 18.5
0.5 SceneSayerODE+ [15] 4.8 5.4 5.4 5.1 6.3 6.9 2.3 3.0 3.0 2.6 3.9 4.4
SceneSayerSDE+ [15] 6.4 7.7 7.9 6.6 8.3 9.0 3.1 3.9 4.0 3.9 5.8 8.4
FDSG (ours) 10.6 13.3 15.5 12.0 15.5 18.8 8.4 11.2 12.6 11.4 17.2 22.5
0.7 SceneSayerODE+ [15] 7.3 7.8 8.0 7.9 8.1 8.4 4.0 4.4 4.5 4.1 5.4 6.5
SceneSayerSDE+ [15] 8.1 8.9 9.3 8.3 9.5 10.6 4.7 5.2 5.4 5.0 6.1 7.4
FDSG (ours) 13.0 15.6 17.8 14.3 18.4 21.5 9.9 12.8 14.1 12.4 18.1 23.1
0.9 SceneSayerODE+ [15] 14.3 17.6 18.0 15.7 18.2 19.4 11.9 13.7 14.4 12.5 14.9 16.8
SceneSayerSDE+ [15] 15.5 18.3 20.6 17.8 20.3 22.8 12.4 14.0 15.3 15.1 16.5 18.7
FDSG (ours) 19.9 23.8 27.6 22.4 27.4 32.5 14.6 18.4 20.7 18.1 24.4 32.9
TABLE IX: Object detection recall and average precision (AP) for SGA and SGF with different observation fraction $\mathcal{F}$.
SGA (IoU $\geq$ 0) SGF (IoU $\geq$ 0.5)
Baseline Oracle Ours Baseline Oracle Ours Baseline Oracle Ours
0.3 77.00 77.05 81.57 37.07 39.23 40.67 9.1 14.0 10.8
0.5 80.11 82.75 83.36 40.33 44.01 43.85 10.5 16.3 12.1
0.7 84.44 89.76 87.38 45.67 51.27 47.87 13.3 23.7 16.7
0.9 86.99 96.01 91.89 55.25 64.95 58.51 18.7 37.5 22.2
We provide further results for Scene Graph Anticipation (SGA) on the Action Genome Scenes (AGS) (Table V), Partially Grounded Action Genome Scenes (PGAGS) (Table VI), and Grounded Action Genome Scenes (GAGS) (Table VII). To provide a more detailed analysis about the anticipation capacity, we evaluate our model under different observation fractions of the video, i.e., $\mathcal{F}=0.3,0.5,0.7,0.9$.
In all three tasks, our model consistently outperforms existing methods in different observation fractions. In AGS, our model significantly outperforms the baselines. In particular, FDSG outperforms SceneSayer [15] across all metrics, achieving up to 50% improvement in mR@10 and mR@50, and averaging a 40% gain across Mean Recall metrics, highlighting our model’s effectiveness in reducing prediction bias. In PGAGS, while SceneSayer [15] marginally outperforms our model on the Recall metrics under the No Constraint setting, FDSG performs notably better than SceneSayer on the Mean Recall metrics, which gives the same indication as AGS results that our model has excellent capability of unbiased scene graph generation. Results in GAGS are consistent with AGS and PGAGS. These results demonstrate the superiority of FDSG in both short-term SGA ($\mathcal{F}=0.9$) and long-term SGA ($\mathcal{F}=0.3$).
In addition, in AGS, the performance gap between SceneSayerSDE and FDSG (e.g., 10.9 in R@20 With Constraint with $\mathcal{F}=0.9$) is clearly greater than that in PGAGS (3.5) and GAGS (3.1). Since ground truth bounding boxes are given in PGAGS and GAGS, this finding provides further support that learning to forecast entity bounding boxes benefits the performance of FDSG.
5.2.3 Scene Graph Forecasting
TABLE X: Ablation study of model components with recall and mean recall metrics. For Dynamic Scene Graph Generation (DSGG) and Scene Graph Forecasting (SGF), the task is SGDET, while for Scene Graph Anticipation (SGA), the task is AGS. For SGA and SGF, the observed fraction $\mathcal{F}=0.5$. Notations for the model components: S - Static SGG Module; T - Temporal Aggregation Module; TDM - Triplet Dynamics Model; LDM - Location Dynamics Model. Best values are highlighted in boldface.
Task # Method Recall@K (R@K)$\uparrow$ Mean Recall@K (mR@K)$\uparrow$
With Constraint No Constraint With Constraint No Constraint
S T TDM LDM 10 20 50 10 20 50 10 20 50 10 20 50
DSGG 1 33.4 41.0 48.1 35.2 45.0 54.0 19.7 25.8 30.9 24.5 38.2 52.0
2 34.6 42.2 49.2 36.5 46.4 55.5 21.5 27.3 33.0 26.2 39.3 50.8
3 34.9 42.4 49.3 36.7 46.6 55.7 21.0 27.2 32.0 26.0 39.5 52.5
4 33.4 41.0 48.1 35.3 45.1 54.2 19.7 25.8 30.9 24.8 38.3 52.2
5 33.3 40.8 47.9 35.2 45.0 53.8 19.7 25.8 30.9 24.6 38.2 51.9
6 35.3 42.9 49.8 37.2 47.4 56.4 22.2 27.8 33.0 27.8 42.0 54.1
SGA 3 28.1 36.5 45.8 29.7 39.9 52.9 14.5 21.5 26.9 17.7 29.9 43.1
4 28.4 36.8 46.3 30.6 40.7 52.7 16.6 21.8 27.2 19.5 31.7 44.5
5 27.2 35.1 44.6 29.0 38.9 50.0 16.2 21.9 28.3 20.8 31.6 44.4
6 28.2 36.5 45.5 29.8 40.1 52.0 18.1 23.5 27.5 23.2 34.3 45.8
SGF 5 9.5 12.2 14.4 10.5 13.9 17.7 6.6 10.1 11.0 8.5 13.6 19.0
6 10.6 13.3 15.5 12.0 15.5 18.8 8.4 11.2 12.6 11.4 17.2 22.5
We report our SGF results in Table VIII. Baselines are adapted from SceneSayerODE/SDE [15], where we use the anticipated relationship representations of SceneSayer LDPU to regress the dynamics of bounding boxes for the corresponding entities starting from the last observed frame. Specifically, we feed the anticipated predicate representations ${z}_{T_{0}+\Delta T}({i,j})$ to two 3-layer MLPs, one for subject bounding box regression and one for object, to update the locations:
$\hat{b}_{T_{0}+\Delta T}(i)=\hat{b}_{T_{0}}(i)+\text{MLP}_{\text{sub}}({z}_{T_{0}+\Delta T}({i,j})),$ (20)
$\hat{b}_{T_{0}+\Delta T}(j)=\hat{b}_{T_{0}}(j)+\text{MLP}_{\text{obj}}({z}_{T_{0}+\Delta T}({i,j})),$ (21)
where both MLPs have a hidden layer dimension of 1,024, with ReLU as the intermediate activation function and SoftPlus as the last activation function. Notebly, in SceneSayer $z$ represents only the predicate, while in our FDSG, $z$ represents the complete triplet. We use the smoothed $L_{1}$ loss for the bounding box anticipation error. Although the baseline models perform well for predicate anticipation, they show limited capacity for entity forecasting. In contrast, our model significantly outperforms them, demonstrating a stronger ability to learn temporal dynamics.
We explicitly evaluate entity forecasting using Object Detection Recall. We evaluate under two settings: (1) SGA, where IoU $\geq 0$, i.e., only evaluates label accuracy; and (2) SGF, where IoU $\geq 0.5$, e.g., includes bounding box quality. We compare against a Baseline using the predicted scene graph from the last observation and an Oracle using the last ground truth. We also report Average Precision (AP) with IoU $\geq 0.5$ in Table IX. Our model clearly outperforms the Baseline, in particular in the SGA setting, showing promising capability to discover novel nodes. Our model sometimes closely trails the Oracle due to minor prediction errors for the last observation, and in the SGF setting due to accumulated error in label and bounding box prediction. However, in the SGF setting, our model still evidently outperforms the Baseline that assumes bounding boxes predicted from the last observation remain static.
FDSG’s superior performance in SGF, together with our improvements in DSGG and SGA (Tables II - VII), highlights the advantages of our forecasting-driven design.